# Optimal Rates for Generalization of Gradient Descent for Deep ReLU Classification

### v1

@@ A.1 Properties of the Initialization
$v^{l}(\mathbf{x}_{i})=\mathbf{H}_{l,0}^{a}(\mathbf{x}_{i})v,\quad a\leq l\leq b.$
and $v^{a-1}(\mathbf{x}_{i})=v$. Note that condition on $h_{0}^{l-1}(\mathbf{x}_{i}),v^{l-1}(\mathbf{x}_{i})$, we take expectation over $\mathbf{w}_{r}^{l}(0)$, applying Lemma 5 implies that
$\mathbb{E}[2\mathbb{I}\{\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle\geq 0\}(\langle\mathbf{w}^{l}_{r}(0),v^{l-1}(\mathbf{x}_{i})\rangle)^{2}]=\|v^{l-1}(\mathbf{x}_{i})\|_{2}^{2}.$
@@
$\|v^{l}(\mathbf{x}_{i})\|^{2}_{2}=\|\mathbf{H}_{l,0}^{a}(\mathbf{x}_{i})v\|_{2}^{2}=\left\|\sqrt{\frac{2}{m}}\Sigma_{0}^{l}(\mathbf{x}_{i})\mathbf{W}^{l}(0)v^{l-1}(\mathbf{x}_{i})\right\|_{2}^{2}$
$=$ $\frac{1}{m}\sum_{r=1}^{m}\mathbb{E}2\mathbb{I}\{\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle\geq 0\}(\langle\mathbf{w}^{l}_{r}(0),v^{l-1}(\mathbf{x}_{i})\rangle)^{2}=\|v^{l-1}(\mathbf{x}_{i})\|_{2}^{2}.$
@@
$\mathbb{P}\left(|\|v^{l}_{j}(\mathbf{x}_{i})\|_{2}^{2}-\|v^{l-1}_{j}(\mathbf{x}_{i})\|_{2}^{2}|\leq\frac{C\|v^{l-1}_{j}(\mathbf{x}_{i})\|_{2}^{2}}{L}|v_{j}^{l-1}(\mathbf{x}_{i})\right)\geq 1-2\exp\left(-Cm\min\left\{\frac{1}{L^{2}},\frac{1}{L}\right\}\right).$
@@
$\mathbb{P}\left(|\|v^{l}_{j}(\mathbf{x}_{i})\|_{2}^{2}-\|v^{l-1}_{j}(\mathbf{x}_{i})\|_{2}^{2}|\leq\frac{C\|v^{l-1}_{j}(\mathbf{x}_{i})\|_{2}^{2}}{L}\right)$
@@
$\|v^{l}_{j}(\mathbf{x}_{i})\|_{2}^{2}\leq\left(1+\frac{C}{L}\right)\|v^{l-1}_{j}(\mathbf{x}_{i})\|_{2}^{2}\leq\left(1+\frac{C}{L}\right)^{L}\|v_{j}\|_{2}^{2}\lesssim 1.$
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v\|_{2}$
$\leq$ $\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v_{j}\|_{2}+\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})(v-v_{j})\|_{2}$
$=$ $\|v_{j}^{b}(\mathbf{x}_{i})\|_{2}+\|v-v_{j}\|_{2}\left\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})\frac{v-v_{j}}{\|v-v_{j}\|_{2}}\right\|_{2}$
$\lesssim$ $1+\|v-v_{j}\|_{2}\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v\|_{2}\leq 1+\frac{1}{2}\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v\|_{2}.$
@@
$\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v\|_{2}\lesssim 1+\frac{1}{2}\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v\|_{2},$
which implies $\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v\|_{2}\lesssim 1$. Hence (23) holds and the proof is completed. ∎
@@
Our proofs are inspired by Lemma A.9 in Zou et al. (2018), which establishes the estimates under the condition $s\gtrsim\log(nL/\delta)$. However, we eliminate this assumption via a more refined analysis.
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})\|_{2}\lesssim L\sqrt{\log m}.$
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v\|_{2}=\left\|\sum_{j=1}^{q}\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v_{j}\right\|_{2}\leq\sum_{j=1}^{q}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v_{j}\|_{2}\leq\sqrt{q\sum_{j=1}^{q}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v_{j}\|_{2}^{2}}$
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})\|_{2}\lesssim L\sqrt{\log m}.$
@@
$\left\|\mathbf{G}^{a}_{b,0}(x_{i})\right\|_{2}\lesssim L\sqrt{\frac{\log m}{m}},\quad i\in[n],1\leq a\leq b\leq L.$ (24)
@@
In Lemma 10, we introduce a useful technique that decomposes the unit vector into sparse components. This approach reduces the covering number from $5^{m}$ to $5^{s}\binom{m}{s}$, making it easier for high-probability bounds to hold. A related method appears in Lemma 7.3 of Allen-Zhu et al. (2019), but their width exhibits polynomial dependence on $n$. In contrast, our analysis achieves polylogarithmic width, substantially relaxing the overparameterization requirement.
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)\mathbf{H}_{b-1,0}^{a}(\mathbf{x}_{i})\Sigma_{2}\right\|_{2}\lesssim\frac{1}{L}.$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)\mathbf{H}_{b-1,0}^{a}(\mathbf{x}_{i})v\right\|_{2}\lesssim\frac{1}{L}.$ (25)
@@
$\mathbb{P}\left(|\|v^{l}_{j}(\mathbf{x}_{i})\|_{2}^{2}-\|v^{l-1}_{j}(\mathbf{x}_{i})\|_{2}^{2}|\leq\frac{C\|v^{l-1}_{j}(\mathbf{x}_{i})\|_{2}^{2}}{L}|v_{j}^{l-1}(\mathbf{x}_{i})\right)\geq 1-2\exp\left(-Cm\min\left\{\frac{1}{L^{2}},\frac{1}{L}\right\}\right).$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)v_{j}^{b-1}(\mathbf{x}_{i})\right\|_{2}^{2}=\sum_{r=1}^{m}d_{r}^{2}\frac{2}{m}(\langle\mathbf{w}^{b}_{r},v_{j}^{b-1}(\mathbf{x}_{i})\rangle)^{2}.$ (26)
Condition on $v^{b-1}_{j}(\mathbf{x}_{i})$, there holds
$\mathbb{E}\frac{2}{m}(\langle\mathbf{w}^{b}_{r}(0),v_{j}^{b-1}(\mathbf{x}_{i})\rangle)^{2}=\frac{2}{m}\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}.$
Let $X_{r}=\frac{2}{m}(\langle\mathbf{w}^{b}_{r}(0),v_{j}^{b-1}(\mathbf{x}_{i})\rangle)^{2}-\frac{2}{m}\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}$ and $d=(d_{1},\cdots,d_{m})$. Then $X_{r}$ are mean-zero sub-exponential random variables, following similar discussions in Lemma 6, we have $\|X_{r}\|_{\phi_{1}}\leq\frac{C}{m}\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}$. Moreover, $\|d\|_{2}^{2}=\sum_{r=1}^{m}d^{2}_{r}=\sum_{r=1}^{m}d_{r}\leq s,\|d\|_{\infty}=1$. Applying Lemma 4, we have
$\mathbb{P}\left(\left|\sum_{r=1}^{m}d_{r}\left(\frac{2}{m}(\langle\mathbf{w}^{b}_{r}(0),v_{j}^{b-1}(\mathbf{x}_{i})\rangle)^{2}-\frac{2}{m}\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}\right)\right|\geq t\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}|v_{j}^{b-1}(\mathbf{x}_{i})\right)$
@@
$\mathbb{P}\left(\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)v_{j}^{b-1}(\mathbf{x}_{i})\right\|_{2}^{2}\leq\left(\frac{2s}{m}+\frac{1}{L^{2}}\right)\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}|v_{j}^{b-1}(\mathbf{x}_{i})\right)$
$\geq$ $\mathbb{P}\left(\sum_{r=1}^{m}d_{r}^{2}\frac{2}{m}(\langle\mathbf{w}^{b}_{r}(0),v_{j}^{b-1}(\mathbf{x}_{i})\rangle)^{2}\leq\sum_{r=1}^{m}\frac{2d_{r}}{m}\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}+\frac{1}{L^{2}}\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}|v_{j}^{b-1}(\mathbf{x}_{i})\right)$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)v_{j}^{b-1}(\mathbf{x}_{i})\right\|_{2}^{2}\leq\left(\frac{2s}{m}+\frac{1}{L^{2}}\right)\|v_{j}^{b-1}(\mathbf{x}_{i})\|_{2}^{2}$
$\leq$ $\left(\frac{2s}{m}+\frac{1}{L^{2}}\right)\left(1+\frac{C}{L}\right)\|v_{j}^{b-2}(\mathbf{x}_{i})\|_{2}^{2}\leq\left(\frac{2s}{m}+\frac{1}{L^{2}}\right)\left(1+\frac{C}{L}\right)^{L}\|v_{j}\|_{2}^{2}\lesssim\frac{1}{L^{2}}.$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)\mathbf{H}_{b-1,0}^{a}(\mathbf{x}_{i})v_{j}\right\|_{2}\lesssim\frac{1}{L}.$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}_{0}(\mathbf{x}_{i})\mathbf{H}_{b-1,0}^{a}(\mathbf{x}_{i})\right\|_{2}\lesssim\sqrt{\log m}.$ (28)

### v2

@@ A.1 Properties of the Initialization
$v^{l}(\mathbf{x}_{i})=\mathbf{H}_{l,0}^{a}(\mathbf{x}^{i})v,\quad a\leq l\leq b.$
and $v^{a-1}(\mathbf{x}^{i})=v$. Note that condition on $h_{0}^{l-1}(\mathbf{x}^{i}),v^{l-1}(\mathbf{x}^{i})$, we take expectation over $\mathbf{w}_{r}^{l}(0)$, applying Lemma 5 implies that
$\mathbb{E}[2\mathbb{I}\{\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle\geq 0\}(\langle\mathbf{w}^{l}_{r}(0),v^{l-1}(\mathbf{x}^{i})\rangle)^{2}]=\|v^{l-1}(\mathbf{x}^{i})\|_{2}^{2}.$
@@
$\|v^{l}(\mathbf{x}^{i})\|^{2}_{2}=\|\mathbf{H}_{l,0}^{a}(\mathbf{x}^{i})v\|_{2}^{2}=\left\|\sqrt{\frac{2}{m}}\Sigma_{0}^{l}(\mathbf{x}^{i})\mathbf{W}^{l}(0)v^{l-1}(\mathbf{x}^{i})\right\|_{2}^{2}$
$=$ $\frac{1}{m}\sum_{r=1}^{m}\mathbb{E}2\mathbb{I}\{\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle\geq 0\}(\langle\mathbf{w}^{l}_{r}(0),v^{l-1}(\mathbf{x}^{i})\rangle)^{2}=\|v^{l-1}(\mathbf{x}^{i})\|_{2}^{2}.$
@@
$\mathbb{P}\left(|\|v^{l}_{j}(\mathbf{x}^{i})\|_{2}^{2}-\|v^{l-1}_{j}(\mathbf{x}^{i})\|_{2}^{2}|\leq\frac{C\|v^{l-1}_{j}(\mathbf{x}^{i})\|_{2}^{2}}{L}|v_{j}^{l-1}(\mathbf{x}^{i})\right)\geq 1-2\exp\left(-Cm\min\left\{\frac{1}{L^{2}},\frac{1}{L}\right\}\right).$
@@
$\mathbb{P}\left(|\|v^{l}_{j}(\mathbf{x}^{i})\|_{2}^{2}-\|v^{l-1}_{j}(\mathbf{x}^{i})\|_{2}^{2}|\leq\frac{C\|v^{l-1}_{j}(\mathbf{x}^{i})\|_{2}^{2}}{L}\right)$
@@
$\|v^{l}_{j}(\mathbf{x}^{i})\|_{2}^{2}\leq\left(1+\frac{C}{L}\right)\|v^{l-1}_{j}(\mathbf{x}^{i})\|_{2}^{2}\leq\left(1+\frac{C}{L}\right)^{L}\|v_{j}\|_{2}^{2}\lesssim 1.$
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v\|_{2}$
$\leq$ $\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v_{j}\|_{2}+\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})(v-v_{j})\|_{2}$
$=$ $\|v_{j}^{b}(\mathbf{x}^{i})\|_{2}+\|v-v_{j}\|_{2}\left\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})\frac{v-v_{j}}{\|v-v_{j}\|_{2}}\right\|_{2}$
$\lesssim$ $1+\|v-v_{j}\|_{2}\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v\|_{2}\leq 1+\frac{1}{2}\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v\|_{2}.$
@@
$\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v\|_{2}\lesssim 1+\frac{1}{2}\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v\|_{2},$
which implies $\sup_{v\in\mathcal{P}}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v\|_{2}\lesssim 1$. Hence (23) holds and the proof is completed. ∎
@@
Our proofs are inspired by Lemma A.9 in zou2018stochasticgradientdescentoptimizes, which establishes the estimates under the condition $s\gtrsim\log(NL/\delta)$. However, we eliminate this assumption via a more refined analysis.
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})\|_{2}\lesssim L\sqrt{\log m}.$
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v\|_{2}=\left\|\sum_{j=1}^{q}\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v_{j}\right\|_{2}\leq\sum_{j=1}^{q}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v_{j}\|_{2}\leq\sqrt{q\sum_{j=1}^{q}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v_{j}\|_{2}^{2}}$
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})\|_{2}\lesssim L\sqrt{\log m}.$
@@
$\left\|\mathbf{G}^{a}_{b,0}(\mathbf{x}^{i})\right\|_{2}\lesssim L\sqrt{\frac{\log m}{m}},\quad i\in[N],1\leq a\leq b\leq L.$ (24)
@@
In Lemma 10, we introduce a useful technique that decomposes the unit vector into sparse components. This approach reduces the covering number from $5^{m}$ to $5^{s}\binom{m}{s}$, making it easier for high-probability bounds to hold. A related method appears in Lemma 7.3 of pmlr-v97-allen-zhu19a, but their width exhibits polynomial dependence on $n$. In contrast, our analysis achieves polylogarithmic width, substantially relaxing the overparameterization requirement.
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)\mathbf{H}_{b-1,0}^{a}(\mathbf{x}^{i})\Sigma_{2}\right\|_{2}\lesssim\frac{1}{L}.$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)\mathbf{H}_{b-1,0}^{a}(\mathbf{x}^{i})v\right\|_{2}\lesssim\frac{1}{L}.$ (25)
@@
$\mathbb{P}\left(|\|v^{l}_{j}(\mathbf{x}^{i})\|_{2}^{2}-\|v^{l-1}_{j}(\mathbf{x}^{i})\|_{2}^{2}|\leq\frac{C\|v^{l-1}_{j}(\mathbf{x}^{i})\|_{2}^{2}}{L}|v_{j}^{l-1}(\mathbf{x}^{i})\right)\geq 1-2\exp\left(-Cm\min\left\{\frac{1}{L^{2}},\frac{1}{L}\right\}\right).$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)v_{j}^{b-1}(\mathbf{x}^{i})\right\|_{2}^{2}=\sum_{r=1}^{m}d_{r}^{2}\frac{2}{m}(\langle\mathbf{w}^{b}_{r},v_{j}^{b-1}(\mathbf{x}^{i})\rangle)^{2}.$ (26)
Condition on $v^{b-1}_{j}(\mathbf{x}^{i})$, there holds
$\mathbb{E}\frac{2}{m}(\langle\mathbf{w}^{b}_{r}(0),v_{j}^{b-1}(\mathbf{x}^{i})\rangle)^{2}=\frac{2}{m}\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}.$
Let $X_{r}=\frac{2}{m}(\langle\mathbf{w}^{b}_{r}(0),v_{j}^{b-1}(\mathbf{x}^{i})\rangle)^{2}-\frac{2}{m}\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}$ and $d=(d_{1},\cdots,d_{m})$. Then $X_{r}$ are mean-zero sub-exponential random variables, following similar discussions in Lemma 6, we have $\|X_{r}\|_{\phi_{1}}\leq\frac{C}{m}\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}$. Moreover, $\|d\|_{2}^{2}=\sum_{r=1}^{m}d^{2}_{r}=\sum_{r=1}^{m}d_{r}\leq s,\|d\|_{\infty}=1$. Applying Lemma 4, we have
$\mathbb{P}\left(\left|\sum_{r=1}^{m}d_{r}\left(\frac{2}{m}(\langle\mathbf{w}^{b}_{r}(0),v_{j}^{b-1}(\mathbf{x}^{i})\rangle)^{2}-\frac{2}{m}\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}\right)\right|\geq t\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}|v_{j}^{b-1}(\mathbf{x}^{i})\right)$
@@
$\mathbb{P}\left(\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)v_{j}^{b-1}(\mathbf{x}^{i})\right\|_{2}^{2}\leq\left(\frac{2s}{m}+\frac{1}{L^{2}}\right)\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}|v_{j}^{b-1}(\mathbf{x}^{i})\right)$
$\geq$ $\mathbb{P}\left(\sum_{r=1}^{m}d_{r}^{2}\frac{2}{m}(\langle\mathbf{w}^{b}_{r}(0),v_{j}^{b-1}(\mathbf{x}^{i})\rangle)^{2}\leq\sum_{r=1}^{m}\frac{2d_{r}}{m}\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}+\frac{1}{L^{2}}\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}|v_{j}^{b-1}(\mathbf{x}^{i})\right)$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)v_{j}^{b-1}(\mathbf{x}^{i})\right\|_{2}^{2}\leq\left(\frac{2s}{m}+\frac{1}{L^{2}}\right)\|v_{j}^{b-1}(\mathbf{x}^{i})\|_{2}^{2}$
$\leq$ $\left(\frac{2s}{m}+\frac{1}{L^{2}}\right)\left(1+\frac{C}{L}\right)\|v_{j}^{b-2}(\mathbf{x}^{i})\|_{2}^{2}\leq\left(\frac{2s}{m}+\frac{1}{L^{2}}\right)\left(1+\frac{C}{L}\right)^{L}\|v_{j}\|_{2}^{2}\lesssim\frac{1}{L^{2}}.$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}(0)\mathbf{H}_{b-1,0}^{a}(\mathbf{x}^{i})v_{j}\right\|_{2}\lesssim\frac{1}{L}.$
@@
$\left\|\Sigma_{1}\sqrt{\frac{2}{m}}\mathbf{W}^{b}_{0}(\mathbf{x}^{i})\mathbf{H}_{b-1,0}^{a}(\mathbf{x}^{i})\right\|_{2}\lesssim\sqrt{\log m}.$ (28)